OpenStudy (anonymous):

Given the functions f(x) = 10x + 25 and g(x) = x + 8, which of the following functions represents f[g(x)] correctly? f[g(x)] = 10x + 33 f[g(x)] = 10x2 + 33 f[g(x)] = 10x + 105 f[g(x)] = 10x2 + 105

OpenStudy (anonymous):

@Nnesha @freckles

OpenStudy (anonymous):

Are you sure those are the answer choices?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you need to replace the x in f(x) with g(x) then simplify f(g(x)) = 10(x + 8) + 25

OpenStudy (anonymous):

So, y=10(10+8)+25?

OpenStudy (anonymous):

f(g(x)) = 10(x + 8) + 25 That's the answer, but you have to simplify/combine like terms. You don't have to substitute a number

OpenStudy (anonymous):

The answer is C. f[g(x)] = 10x + 105

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

lmao. you still don't know how to do it :/

OpenStudy (anonymous):

May you explain it to me then?

OpenStudy (anonymous):

f(g(x)) means to literally put g(x) inside of f(x).